# Optimizing the Network Topology of a Linear Reservoir Computer

### v1

@@ Abstract
Machine learning has become a fundamental approach for modeling, prediction, and control, enabling systems to learn from data and perform complex tasks. Reservoir computing is a machine learning tool that leverages high-dimensional dynamical systems to efficiently process temporal data for prediction and observation tasks. Traditionally, the connectivity of a reservoir computer (RC) is generated at random, lacking a principled design. Here, we focus on optimizing the topology of a linear RC to improve its performance and interpretability, which we achieve by decoupling the RC dynamics into a number of independent modes. We then proceed to optimize each one of these modes to perform a given task, which corresponds to selecting an optimal RC connectivity in terms of a given set of eigenvalues of the RC adjacency matrix. Simulations on networks of varying sizes show that the optimized RC significantly outperforms randomly constructed reservoirs in both the training and testing phases and also often surpasses nonlinear reservoirs of comparable size. This approach provides both practical performance advantages and theoretical guidelines for designing efficient, task-specific, and analytically transparent RC architectures.
@@ 1 Introduction
Machine learning has emerged as a key approach for modeling and control, allowing systems to extract patterns from data and perform complex tasks with high accuracy Zhao et al. (2024); Bensoussan et al. (2022); Hashemi et al. (2023). However, training in many machine learning can be computationally demanding. Reservoir computing, a machine learning tool within the broader class of recurrent neural networks, addresses this challenge by keeping the internal reservoir connectivity fixed and training only the readout layer Jaeger (2001). This allows to perform training in terms of a simple regression operation, as opposed to the more computationally intensive back-propagation. RC has shown to be highly successful in processing temporal and sequential data, especially in prediction and observation tasks McCreesh and Cortés (2024); Shen et al. (2025); Pathak et al. (2018); Jordanou et al. (2022).
@@
In typical implementations, reservoir computers employ nonlinear node dynamics, and their network topologies are chosen at random Yan et al. (2024). Although this random construction often yields satisfactory results in many tasks, it lacks a principled foundation and is not explicitly optimized to perform given tasks of interest Dale et al. (2021). Furthermore, the presence of nonlinearities complicates the analysis of the internal dynamics, making it difficult to characterize key properties such as memory capacity, stability, and input-output relationships. Metzner et al. (2025); Sugiura et al. (2024). Consequently, nonlinear RC often function as black boxes, providing limited insight into how information is processed or stored. In contrast, linear reservoir computing (LRC) significantly improves interpretability and analytical tractability Gauthier et al. (2021). Although linear reservoirs may exhibit lower performance in general, they can be highly effective for tasks with well-understood linear or frequency-driven structures, offering a valuable platform for developing theory-guided, efficient, and generalizable RC systems. Our approach in this paper adopts linear reservoir dynamics, which allows systematic optimization of the network topology to enhance performance on specific tasks of interest.
@@
In what follows, we consider an LRC performing an observation task, in which the reservoir is trained to reconstruct an output signal from knowledge of an input signal Lu et al. (2017). Our objective is to optimize the network topology of the LRC, based on knowledge of the spectral characteristics of the input and output signals. Specifically, we formulate an optimization problem to determine the optimal set of eigenvalues of the matrix that provides the connectivity between reservoir nodes, which we show leads to a drastic reduction of the training error. By leveraging frequency-domain insights, we identify reservoir topologies that maximize performance and interpretability. Our results show that the optimized LRC not only outperforms randomly constructed LRC but also achieves better performance than nonlinear RC of comparable size.
@@ 2 System Model and Reservoir Dynamics
We take two signals which we know to be dynamically related to one another, $\hat{u}(t)$ and $\hat{y}(t)$. Fundamentally, our goal is to ‘observe’ or ‘learn’ $\hat{y}(t)$ from knowledge of $\hat{u}(t)$. To this end, we extract a set of $K$ dominant frequencies that are simultaneously present in both $\hat{y}(t)$ and $\hat{u}(t)$, and construct the signals $y(t)$ and $u(t)$ that only contain those frequencies Butschek et al. (2022). Hence, $u(t)$ and $y(t)$ can be expressed as:
@@
To this end, we introduce a linear reservoir computer, which obeys the dynamics,
@@
where $\mathbf{r}(t)=[r_{1}(t),r_{2}(t),...,r_{N}(t)]$ is the vector of nodes’ readout, $\gamma$ is a positive constant, the symmetric matrix $A$ is the reservoir’s adjacency matrix, i.e., $A_{ij}=A_{ji}$ describes the strength of the coupling between node $j$ to node $i$, the mask vector $\mathbf{d}=[d_{1},d_{2},...,d_{N}]$, $N$ is the number of nodes, and $u(t)$ is the input signal. We also assume that the matrix $A$ is Hurwitz, and therefore the reservoir dynamics (3) is stable.
To compute the reservoir modes, the training signal is discretized using a time step $\tau$. The reservoir produces an estimate of the training signal, the fit signal $\mathbf{\check{y}}$, in terms of a linear combination of the readout matrix $\Omega_{r}$ in the coefficient vector, $\boldsymbol{\kappa}$
$\Omega_{r}\boldsymbol{\kappa}=\mathbf{\check{y}},$ (4)
where
$\Omega_{r}=\left(\begin{array}[]{ccccc}r_{1}(\tau)&r_{2}(\tau)&...&r_{N}(\tau)&1\\ r_{1}(2\tau)&r_{2}(2\tau)&...&r_{N}(2\tau)&1\\ \vdots&\vdots&\vdots&\vdots&\\ r_{1}(T\tau)&r_{2}(T\tau)&...&r_{N}(T\tau)&1\\ \end{array}\right)$ (5)
and $T$ is the number of time steps. We wish to choose the coefficients of the vector $\boldsymbol{\kappa}$ that minimizes the error $\epsilon_{r}$,
@@
where $||.||$ denotes the $L_{2}$ norm. We compute $\boldsymbol{\kappa}^{*}=\Omega_{r}^{\dagger}\mathbf{y}$ the optimal vector of the output weights that minimizes $\epsilon_{r}$ in Eq. (6), where ${\Omega_{r}}^{\dagger}=\left(\Omega_{r}^{T}\Omega_{r}+\beta I\right)^{-1}\Omega_{r}^{T}$ and $\beta\geq 0$ is the ridge regression parameter Wyffels et al. (2008).
To facilitate optimization of the linear reservoir, we employ modal decomposition to transform the reservoir dynamics into independent modes that can be individually tuned. To do so, since the matrix $A$ is symmetric, it can be diagonalized as $A=V\Lambda V^{T}$, where $V$ is an orthogonal matrix whose columns are the eigenvectors of $A$ and $\Lambda$ is a diagonal matrix with the real eigenvalues of $A$ along its main diagonal. We pre-multiply both sides of Eq. (3) by $V^{T}$, and by defining $\mathbf{q}(t)=V^{T}\mathbf{r}(t)$ and $\mathbf{c}=V^{T}\mathbf{d}$, the system transforms into the following form,
@@
where $\mathbf{q}$ is the vector of the decoupled reservoir readouts, which breaks up into a set of $N$ independent modes,
@@
where $\lambda_{i}$ are the eigenvalues of the reservoir’s adjacency matrix $A$.
Equations (7) and (8) provide a modal decomposition for the reservoir dynamics. The implications of the modal decomposition of an RC are significant and are illustrated in Fig. 1. Panel (a) of this figure shows a coupled reservoir, where nodes are interconnected, whereas panel (b) illustrates a decoupled reservoir computer, in which the reservoir is decomposed into independent modes without node-to-node connections. In particular, the fact that an RC can be decomposed into a number of independent modes (see Fig. 1) has important implications and arguably several advantages in terms of RC robustness, scalability, repurposability, and interoperability. Another fundamental advantage is the possibility of more easily optimizing the structure of a reservoir, as it reduces the parameter space over which the optimization is performed from $N^{2}$ entries of the matrix $A$ to only $N$ eigenvalues. This is specifically what we do in this paper.
@@
The decoupled reservoir is easier to handle since it consists of independent rather than dependent modes. An important step is to understand the error relationship between the coupled reservoir of Eqs. (3) and the decoupled reservoir of Eqs. (7). This relationship is formally established in Theorem 1.
@@
Let a coupled linear reservoir and a decoupled linear reservoir be described by Eqs. (3) and (7), respectively. Then, the best-fit signals for the two problems are identical.
@@
The detailed proof is provided in Appendix I. ∎
Since the decoupled reservoir is simpler to work with and it yields the same training and testing errors as the coupled reservoir, in the remainder of the paper we focus on the decoupled reservoir model. Notwithstanding the benefits of the decoupled reservoir, the dimension of the reservoir mode matrix remains large, as it scales with the number of time steps. This dimension can be significantly reduced by transforming the problem into the frequency domain, where the matrix size depends only on the number of input frequencies, which is considerably smaller than the number of time steps. To this end, we reformulate the problem in the frequency domain by applying the Laplace transform to each mode and expressing it through its corresponding transfer function. The transfer function of each mode in Eq. (9) can be expressed as follows.
@@ 2.1 Generic Relation Between the Error In The Time and the Frequency Domain
The reservoir state matrix associated with Eq. (11) is:
$\Omega_{q}=\begin{bmatrix}q_{1}(1)&q_{2}(1)&\dots&q_{N}(1)\\ q_{1}(2)&q_{2}(2)&\dots&q_{N}(2)\\ \vdots&\vdots&\ddots&\vdots\\ q_{1}(T)&q_{2}(T)&\dots&q_{N}(T)\end{bmatrix}$ (13)
@@
The output of the reservoir computer is constructed as a linear combination of the internal states, referred to as the readouts. By taking the Laplace transform of Eq. ( 11), we obtain,
@@
At the same time, the Laplace transform of the output signal of Eq. ( 2) is as follows,
@@
where $\tilde{\Omega}_{q}$ is given by Eq. (26) and
@@
One of the benefits of computing $\boldsymbol{\tilde{\kappa}}^{*}$ as opposed to computing $\boldsymbol{\kappa}^{*}$ is that the matrix $\tilde{\Omega}_{q}$ in Eq. (26) has a number of rows equal to 2KK, where $K$ is the number of frequencies, which is typically lower than the number of rows $T$ of the matrix ${\Omega}_{q}$ in Eq. (13). This results in lower computational cost, a simplified interpretation and analysis, and more efficient calculations. While especially useful for periodic or narrowband inputs, understanding the relationship between errors in the time and frequency domains remains essential, due to the inherently temporal nature of learning in reservoirs. Theorem 2 explicitly establishes this relationship.
@@
Let $\Omega_{q}\in\mathbb{R}^{T\times N}$ be the reservoir state matrix of a linear decoupled reservoir computer, and let $\boldsymbol{y}\in\mathbb{R}^{T}$ be the corresponding target output signal in the time domain. Define the time-domain readout weights $\boldsymbol{\kappa}\in\mathbb{R}^{N}$ as:
@@
where $\Omega_{q}^{\dagger}={\left(\Omega_{q}^{T}\Omega_{q}+\beta I\right)^{-1}\Omega_{q}^{T}}$. Let $\widetilde{\Omega}_{q}$ and $\tilde{\boldsymbol{y}}$ be the discrete Fourier transforms (DFT) of $\Omega_{q}$ and $\boldsymbol{y}$, respectively. Then, the frequency-domain readout weights,
@@
are equal to the time-domain readout weights, i.e.,
@@
The detailed proof is provided in Appendix II. ∎
@@
$\tilde{\epsilon}_{q}\approx\epsilon_{q}\approx\epsilon_{r}\approx\epsilon$ (27)
@@
While $\epsilon_{r}$ and its frequency-domain counterpart are useful for theoretical analysis, in our numerical experiments we use the Normalized Root Mean Squared Error (NRMSE) to compare the performance of different reservoir architectures.
$\mathrm{NRMSE}=\frac{\|\check{y}-y\|}{\|y\|},$ (28)
where $\check{y}$ is the approximation of the target signal, $y$ is the true (measured or reference) signal.

### v2

@@ Abstract
Machine learning has become a fundamental approach for modeling, prediction, and control, enabling systems to learn from data and perform complex tasks. Reservoir computing is a machine learning tool that leverages high-dimensional dynamical systems to efficiently process temporal data for prediction and observation tasks. Traditionally, the connectivity of the network that underlies a reservoir computer (RC) is generated randomly, lacking a principled design. Here, we focus on optimizing the connectivity of a linear RC to improve its performance and interpretability, which we achieve by decoupling the RC dynamics into a number of independent modes. We then proceed to optimize each one of these modes to perform a given task, which corresponds to selecting an optimal RC connectivity in terms of a given set of eigenvalues of the RC adjacency matrix. Simulations on networks of varying sizes show that the optimized RC significantly outperforms randomly constructed reservoirs in both training and testing phases and often surpasses nonlinear reservoirs of comparable size. This approach provides both practical performance advantages and theoretical guidelines for designing efficient, task-specific, and analytically transparent RC architectures.
@@ 1 Introduction
Machine learning has emerged as a key approach for modeling and control, allowing systems to extract patterns from data and perform complex tasks with high accuracy Zhao_2025; BENSOUSSAN2022531; Hashemi2023. However, training in machine learning can be computationally demanding. Reservoir computing (RC), a machine learning tool within the broader class of recurrent neural networks, addresses this challenge by keeping the internal reservoir connectivity fixed and training only the readout layer jaeger2001echo. This allows training to be performed as a simple regression operation, as opposed to the more computationally intensive back-propagation. RC has been shown to be highly successful in processing temporal and sequential data, especially in prediction and observation tasks 10659224; 10816480; PhysRevLett.120.024102; 9664461; del2021reservoir; nathe2023reservoir.
In typical implementations, reservoir computers employ nonlinear node dynamics; moreover, the network topology, i.e., the node-to-node connectivity of the reservoir, is chosen randomlyYan2024.
Although this random construction often yields satisfactory results in many tasks, it lacks a principled foundation and is not explicitly optimized to perform specific tasks of interest Dale2021. This points out to an inherent trade-off between performance and repurposability, i.e., one can either use a reservoir with a random topology and expect it to perform different tasks well; or can pick a specific topology that is optimized to perform a specific task, but not others. In the rest of this paper, we focus on optimizing the structure of a RC to perform specific tasks of interest.
@@
Furthermore, the presence of nonlinearities complicates the analysis of the RC internal dynamics, making it difficult to characterize key properties such as memory capacity, stability, and input-output relationships 10.1162/neco_a_01770; 10221724. Consequently, nonlinear RCs often function as black boxes, providing limited insight into how information is processed or stored. In contrast, linear reservoir computing (LRC) significantly improves interpretability and analytical tractability Gauthier2021. Although linear reservoirs may exhibit lower performance in general, they can be highly effective for tasks with well-understood linear or frequency-driven structures, offering a valuable platform for developing theory-guided, efficient, and generalizable RC systems. Our approach in this paper adopts linear reservoir dynamics, which we will show allows systematic optimization of the network topology to enhance performance on specific tasks of interest.
@@
In what follows, we consider an LRC performing an observation task, in which the reservoir is trained to reconstruct an output signal from knowledge of an input signal 10.1063/1.4979665. Our objective is to optimize the network topology of the LRC, based on knowledge of the spectral characteristics of the input and output signals. Specifically, we formulate an optimization problem to determine the optimal set of eigenvalues of the adjacency matrix that provides the connectivity between reservoir nodes, which we show leads to a drastic reduction of the training error. By leveraging frequency-domain insights, we identify reservoir topologies that maximize performance and interpretability. Our results show that the optimized LRC not only outperforms randomly constructed LRC but also often achieves better performance than nonlinear RCs of comparable size.
@@ 2 System Model and Reservoir Dynamics
We take two signals which we know to be dynamically related to one another, $\hat{u}(t)$ and $\hat{y}(t)$. Fundamentally, our goal is to ‘observe’ or ‘learn’ $\hat{y}(t)$ from knowledge of $\hat{u}(t)$. To this end, we extract a set of $K$ dominant frequencies that are simultaneously present in both $\hat{y}(t)$ and $\hat{u}(t)$, and construct the signals $y(t)$ and $u(t)$ that only contain those frequencies Butschek:22. Hence, $u(t)$ and $y(t)$ can be expressed as:
@@
To this end, we introduce a linear reservoir computer with $N$ nodes, which obeys the dynamics,
@@
where $\mathbf{r}(t)=[r_{1}(t),r_{2}(t),...,r_{N}(t)]$ is the $N$-dimensional vector of the nodes’ readout, $\gamma$ is a positive constant, and $A$ is the square $N$-dimensional reservoir’s adjacency matrix, i.e., $A_{ij}$ describes the strength of the coupling from node $j$ to node $i$, the mask vector $\mathbf{d}=[d_{1},d_{2},...,d_{N}]$, and $u(t)$ is the input signal. We sometimes simply refer to the number of nodes $N$ as the reservoir ‘size’. We introduce two assumptions: (i) the matrix $A$ is diagonalizable and (ii) the matrix (AI)(A-I) is Hurwitz. As a result of (ii) and under the generic assumption of bounded inputs $\mathbf{u}$, the reservoir dynamics (3) is stable.
The readouts and the training signal are discretized with a time step $\tau$ and collected in the readout matrix $\Omega_{r}$ and in the training vector $\mathbf{y}$, respectively,
$\Omega_{r}=\left(\begin{array}[]{ccccc}r_{1}(\tau)&r_{2}(\tau)&...&r_{N}(\tau)&1\\ r_{1}(2\tau)&r_{2}(2\tau)&...&r_{N}(2\tau)&1\\ \vdots&\vdots&\vdots&\vdots&\\ r_{1}(T\tau)&r_{2}(T\tau)&...&r_{N}(T\tau)&1\\ \end{array}\right),\quad\quad\mathbf{y}=\left[\begin{array}[]{c}y(\tau)\\ y(2\tau)\\ ...\\ y(T\tau)\end{array}\right],$ (4)
where $T$ is the number of time steps. Throughout this paper $\tau$ is fixed and equal $0.01$; therefore, the duration of the training phase is completely determined by the number of steps $T$.
@@
[…]
Once we have learned the optimal output weights $\boldsymbol{\kappa}^{*}$, we can also compute the testing error lukosevicius2009reservoir,
[…]
where $\check{\Omega}_{r}$ and $\check{\mathbf{y}}$ are computed in the exact same way as ${\Omega}_{r}$ and $\mathbf{y}$}, but over a different interval of time. In the rest of this paper, the duration of the testing phase is taken to be one third of that of the training phase, i.e, the number of samples during the testing phase is equal to $\lceil T/3\rceil$ and the symbol $\lceil\cdot\rceil$ indicates ceil-rounding to the closest integer.
To facilitate optimization of the linear reservoir, here we introduce a modal decomposition to transform the reservoir dynamics into independent modes that can be individually tuned. Since the matrix $A$ is diagonalizable, it can be rewritten in the form $A=V\Lambda V^{-1}$, where $V$ is the matrix whose columns are the eigenvectors of $A$ and $\Lambda$ is a diagonal matrix with the possibly complex eigenvalues of $A$ along its main diagonal. We pre-multiply both sides of Eq. (3) by $V^{-1}$, and by defining the vectors $\mathbf{q}(t)=V^{-1}\mathbf{r}(t)$ and $\mathbf{c}=V^{-1}\mathbf{d}$, the system transforms into the following form,
@@
where $\mathbf{q}(t)$ is the $N$-dimensional vector of the decoupled reservoir readouts. Equation (7) breaks up into a set of $N$ independent modes,
@@
where $\lambda_{i}$ are the eigenvalues of the reservoir’s adjacency matrix $A$. In general, both the modes $q_{i}(t)$ and the eigenvalues $\lambda_{i}$ can be complex, but in the case in which $A$ is symmetric, they are real.
Equations (7) and (8) provide a modal decomposition for the reservoir dynamics.
The implications of the modal decomposition of an RC are significant and are illustrated in Fig. 1. Panel (a) of this figure shows a coupled reservoir, where nodes are interconnected, whereas panel (b) illustrates a decoupled reservoir computer, in which the reservoir is decomposed into independent modes without node-to-node connections. In particular, the fact that an RC can be decomposed into a number of independent modes (see Fig. 1) has important implications and arguably several advantages in terms of RC robustness, scalability, and modularity. Next we briefly comment on each one of these advantages. Robustness arises from the fact that intentional or accidental removal of a single node (or of a set of the nodes) will not propagate to the others. Scalability is achieved since increasing the dimension of the reservoir simply corresponds to adding more individual nodes. This may have a considerable advantage in experimental setting, as it removes the need to wire different nodes in a complicated network structure. Finally, modularity is achieved when a complex structure is replaced by smaller, manageable, and interchangeable components. Another fundamental advantage is the possibility of more easily optimizing the structure of a reservoir, as it reduces the parameter space over which the optimization is performed from $N^{2}$ entries of the matrix $A$ to only $N$ eigenvalues. This is specifically what we do in this paper.
@@
Similar to what done before, we can construct the readout matrix for the decoupled reservoir,
$\Omega_{q}=\left(\begin{array}[]{ccccc}q_{1}(\tau)&q_{2}(\tau)&...&q_{N}(\tau)&1\\ q_{1}(2\tau)&q_{2}(2\tau)&...&q_{N}(2\tau)&1\\ \vdots&\vdots&\vdots&\vdots&\\ q_{1}(T\tau)&q_{2}(T\tau)&...&q_{N}(T\tau)&1\\ \end{array}\right)$ (10)
and compute the error
[…]
where $\boldsymbol{\kappa}_{q}^{*}=\Omega_{q}^{\dagger}\mathbf{y}$.
The decoupled reservoir is easier to handle since it consists of independent rather than dependent modes. An important step is to understand the relationship between the error $\epsilon_{r}$ for the coupled reservoir of Eqs. (3) and the error $\epsilon_{q}$ for the decoupled reservoir of Eqs. (7). This relationship is formally established in Theorem 1.
@@
Assume the matrix $A$ is diagonalizable. Let a coupled linear reservoir and a decoupled linear reservoir be described by Eqs. (3) and (7), respectively. Then, if either AA is normal or $\beta=0$, we have that $\epsilon_{q}=\epsilon_{r}$.
@@
We construct the matrix
$W=\begin{bmatrix}V&\mathbf{0}_{N\times 1}\\ \mathbf{0}_{1\times N}&1\end{bmatrix}$,
where $V$ is a generic invertible matrix $V$. The inverse of $W$ is equal to
$W^{-1}=\begin{bmatrix}V^{-1}&\mathbf{0}_{N\times 1}\\ \mathbf{0}_{1\times N}&1\end{bmatrix}$.
We recall that $\mathbf{q}(t)=V^{-1}\mathbf{r}(t)$. Thus we have $\Omega_{q}=\Omega_{r}W^{-1}$ and $\Omega_{q}^{T}={W^{-1}}^{T}\Omega_{r}^{T}$. By plugging in the expression for the optimal $\boldsymbol{\kappa}_{q}^{*}=\Omega_{q}^{\dagger}\textbf{y}$ in Eq. (11) of the main manuscript, we obtain
[…]
where the operation $\Omega_{r}^{\dagger\beta,W}=(\Omega_{r}^{T}\Omega_{r}+\beta W^{T}W)^{-1}\Omega_{r}^{T}$ is similar to a ridge regression but with the ridge regression parameter $\beta$ replaced by the term $\beta W^{T}W$, which also depends on the particular choice of the matrix $W$.
In general we can conclude that $\epsilon_{q}=\epsilon_{r}^{W}$, where the error $\epsilon_{r}^{W}$ is computed by performing the alternative regression that uses $\Omega_{r}^{\dagger W}$ in place of $\Omega_{r}^{\dagger}$. We also see that there are two cases under which $\epsilon_{q}=\epsilon_{r}$: (i) when the matrix $A$ is normal, which implies $W^{T}W=I$ and (ii) when $\beta=0$. Also, for small $\beta$, $\epsilon_{q}\simeq\epsilon_{r}$.
In the rest of this paper we focus on the decoupled reservoir of Eqs. (7) since it is simpler to work with and it yields the same training and testing errors as the coupled reservoir (Theorem 1.) Notwithstanding the benefits of the decoupled reservoir, the dimension of the readout matrix remains large, as it scales with the number of time steps. This dimension can be significantly reduced by transforming the problem into the frequency domain, where the matrix size depends only on the number of input frequencies $K$, which in practice is considerably smaller than the number of time steps $T$. To this end, we reformulate the problem in the frequency domain by applying the Laplace transform to each mode and expressing it through its corresponding transfer function.
In what follows, proceed under the assumption that the matrix $A$ has a real spectrum and that the modes $q_{i}(t)$ are real. The transfer function of each mode in Eq. (9) can be expressed as follows,
@@ 2.1 Generic Relation Between the Error In The Time and the Frequency Domain
The output of the reservoir computer is constructed as a linear combination of the internal states, referred to as the readouts.
By taking the Laplace transform of Eq. ( 14), we obtain,
@@
At the same time, the Laplace transform of the output signal in Eq. ( 2) is,
@@
where $\tilde{\Omega}_{q}$ is given by
and
@@
The main benefits of computing $\boldsymbol{\tilde{\kappa}}^{*}$, as opposed to computing $\boldsymbol{\kappa}^{*}$, is that the matrix $\tilde{\Omega}_{q}$ in Eq. (20) has a number of rows equal to 2KK, where $K$ is the number of frequencies, which is typically much lower than the number of rows $T$ of the matrix ${\Omega}_{q}$ in Eq. (10). This results in lower computational cost, a simplified interpretation and analysis, and more efficient calculations. While especially useful for periodic or narrowband inputs, an essential step is to understand the relationship between errors in the time and frequency domains, due to the inherently temporal nature of learning in reservoirs. Theorem 2 explicitly establishes this relationship.
@@
Let $\Omega_{q}\in\mathbb{R}^{T\times N}$ be the reservoir readout matrix of a linear decoupled reservoir computer, and let $\boldsymbol{y}\in\mathbb{R}^{T}$ be the corresponding training signal in the time domain. Define the time-domain readout weights $\boldsymbol{\kappa}\in\mathbb{R}^{N}$ as:
@@
where as specified before, $\Omega_{q}^{\dagger}={\left(\Omega_{q}^{T}\Omega_{q}+\beta I\right)^{-1}\Omega_{q}^{T}}$. Let $\widetilde{\Omega}_{q}$ and $\tilde{\boldsymbol{y}}$ be the discrete Fourier transforms (DFT) of $\Omega_{q}$ and $\boldsymbol{y}$, respectively. Then, the frequency-domain readout weights,
@@
are approximately equal to the time-domain readout weights, i.e.,
@@
The detailed proof is provided in Section S2 of the supplementary material. The errors are only approximate as there is always a discretization error. This equivalence improves with a closer discretization approximation. ∎
@@
$\tilde{\epsilon}_{q}\approx\epsilon_{q}=\epsilon_{r}$ (22)
@@
While $\epsilon_{r},\epsilon_{q}$, and $\tilde{\epsilon}_{q}$ are useful for theoretical analysis, in our numerical experiments we use the Normalized Root Mean Squared Error (NRMSE) to compare the performance of different reservoir architectures.
$\mathrm{NRMSE}=\frac{\|\textbf{y}-\textbf{y}^{\text{fit}}\|}{\|\textbf{y}\|},$ (23)
where $\textbf{y}^{\text{fit}}$ is the best-fit approximation of the target signal and y is the true (measured or reference) signal.